ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

1. Logic

1991 MS Classification: 54, 18 Preuß, Gerhard, Free University of Berlin, Germany Convenient Topology - A New Branch of Topology Convenient Topology consists essentially in the study of semiuniform convergence spaces. These ones include not only convergence structures such as topological structures and limit space structures, but also uniform convergence structures such as uniform structures and uniform limit space structures. They are suitable for studying continuity, Cauchy continuity and uniform continuity as well as convergence structures in function spaces, namely simple convergence, continuous convergence and uniform convergence. Several results are presented which cannot be obtained by using topological or uniform spaces respectively. Reference: G. Preuß, Convenient Topology, Math. Japonica 47, 1771-182